A man deposited $5000 in a bank paying 6% simple interest. How many years will it take to earn $6000 in interest? Use I=Prt

Answer by rfer(16322)

I=Prt
t=I/Pr
t=6000/5000*.06
t=6000/300
t=20 yrs